OpenStudy (anonymous):

Which value of c from the set {4, 5, 6, 7, 8} makes this equation true? 2c + 9 = 25

OpenStudy (anonymous):

two options. 1. solve the equation for c. 2. plug in each value of the set for c. The one that gives you 25 for the left side is the answer.

OpenStudy (anonymous):

how do you want to approach this?

OpenStudy (anonymous):

4 5 6 7 8 Are the choices.

OpenStudy (anonymous):

yeah, I got that from above. how do you want to go about solving the problem?

OpenStudy (anonymous):

I dont know.

OpenStudy (anonymous):

I gave you two different approaches above...

OpenStudy (anonymous):

Ohh.

OpenStudy (anonymous):

The second one?

OpenStudy (anonymous):

ok. The first number is 4, so start by replacing c with 4 in the equation. 2c + 9 = 25 2*4 + 9 = 25 8 +9 = 25 17 = 25 → this is false, so 4 isn't the answer.

OpenStudy (anonymous):

Do this for each number in the set

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

did you get it?